OpenStudy (anonymous):

The peremiter of a rectangular swimming pool is 160 feet.Jan wants to swim across the diagonal of the pool. How many times will she have to cross this way to swim at least 200 feet.

jhonyy9 (jhonyy9):



jhonyy9 (jhonyy9):

using pythagora diagonal squared = 3600 +400 = 4000 length of the diagonal = sqrt(4000) =20sqrt10 feet